1. 16 – 8 and 1/2= 8 and 1/2 true or false?

2. The largest fraction listed below is:

a. 1/2.

b. 1/4.

c. 1/3.

d. 1/5.

3. 35.6 – 0.99 =

4. Lanoxin 0.125 mg is to be given. The nurse converts the dose to how many grams?

a. 1.250

b. 0.000125

c. 0.00125

d. 1250

5. Choose the correct abbreviation for both ears.

a. os

b. ad

c. au

d. ou

6. Choose the correct abbreviation for bedtime

a. noc

b. hos

c. bt

d. hs

7. 4:6::2:X

complete the ratio. X =______

8. A patient is prescribed potassium 15 mEq in an oral solution. How many ounces is the correct dose to be administered to the patient from a solution of potassium with a strength of 20 mEq in 8 ounces?

_____ ounces

9. A patient is prescribed 4 mg of morphine IV. The drug on hand is morphine 10 mg in a 1 mL syringe. How many milliliters is the correct dose?

_____ mL

10. The doctor orders 5mg of Versed IV push and 75mcg of Fentanyl IV push. The versed is 5mg/5ml and the Fentanyl is 100mcg/2ml. How many ml’s of Versed and Fentanyl will the patient receive?

a. 5ml Versed/ 0.75ml Fentanyl

b. 2.5ml Versed/ 1.25ml Fentanyl

c. 5ml Versed/ 1.5ml Fentanyl

d. 5ml Versed/ 2ml Fentanyl
Answer #1

Answer- 4) b (0.000125)

Answer- 5) c (au)

Answer- 6)  d (hs)

Answer -8) given information is-prescribed potassium - 15 mEq

Strength of potassium - 20 mEq in 8 ounces

to find - how many ounces is the correct dose to be administered

find- x = D/ A × Q , where D= desired dose or doctors prescription

A= strength of medicine or strength printed on medicine level

  Q= unit in which printed strength present

X = 15 mEq/ 20mEq × 8 ounces

= 6 ounces ( answer)

Answer- 9) given information , prescribed morphine 4 mg I/V

Strength of morphine given is 10 mg in a 1 ml syringe

to find-  how many millilitre is the correct dose

find- x= D/ A× Q

= 4 mg / 10mg × 1ml

= .4ml ( answer)

Answer 10) given information,  Prescribed versed 5 mg IV

Prescribed fentanyl 75 mcg IV

Strength of versed = 5 mg/ 5 ml

Strength of fentanyl= 100 mcg/2ml

to find- how many ml's of versed and fentanyl will patient receive

Find- for versed = X = D/ A× Q

  = 5mg/ 5mg × 5ml

= 5 ml versed

For fentanyl = X= D/ A× Q

= 75mcg/100much×2 ml

= 1.5ml fentanyl

Answer is (c) 5ml versed/1.5 ml fentanyl